OpenStudy (anonymous):

what is the simplified form of 9x^2-25/3x-5

OpenStudy (anonymous):

Factor the top

OpenStudy (anonymous):

what did u get

OpenStudy (anonymous):

cause i got 3x+5, restrictions x not equal to -5/3

OpenStudy (anonymous):

The restricion on the domain is wrong.

OpenStudy (anonymous):

so what did u get

OpenStudy (anonymous):

The reduced form is correst but the restricion is that x =/= 5/3

OpenStudy (anonymous):

why is it negative

OpenStudy (anonymous):

What is negative?

OpenStudy (anonymous):

why isnt it negative

OpenStudy (anonymous):

Because the point restricted from your domain is the point that makes the denominator zero 3x-5=0 3x=5 x=5/3

OpenStudy (anonymous):

its 3x+5=0

OpenStudy (anonymous):

That's now what your wrote up there

OpenStudy (anonymous):

(3x-5)(3x+5)/3x-5; 3x-5 would cancel out leaving 3x+5

OpenStudy (anonymous):

Yes but the denominator of the original expression is 3x-5. The point that makes that zero is the one excluded from the domain of the simplified one.

OpenStudy (anonymous):

ok i have another question

OpenStudy (anonymous):

wouldnt they cancel out